OpenStudy (anonymous):

What is the line of symmetry for the parabola whose equation is y = -x2 + x + 3?

OpenStudy (anonymous):

x=-1 x=-1/2 x=1/2

OpenStudy (anonymous):

complete the square on R.H.S

OpenStudy (anonymous):

\[y=-\left( x^2-x+\frac{ 1 }{ 4}-\frac{ 1 }{ 4 } \right)+3\]

OpenStudy (anonymous):

\[y=-\left( x-\frac{ 1 }{ 2 } \right)^2+\frac{ 1 }{ 4 }+3\] \[\left( x-\frac{ 1 }{ 2 } \right)^2=-\left( y+\frac{ 13 }{ 4 } \right)\] can you find your solution.

OpenStudy (anonymous):

yes thanks